# PMGDA: A Preference-based Multiple Gradient Descent Algorithm

### 14 Feb 2024 v1

@@ IV Experiments
We conducted experiments on three types of problems: synthetic problems, multitask learning problems (fairness classification), and multiobjective reinforcement learning problems. For synthetic problems, the decision vector ${\bm{\theta}}$ is a 30-D vector. For the other two problems, the decision variables are the parameters of neural networks (this number is around 5000-10000).
@@
    The cross angle indicator $\vartheta=\text{arccos}\frac{{\bm{L}}\cdot{\bm{\lambda}}}{\left\|{\bm{L}}\right\|\left\|{\bm{\lambda}}\right\|}$. A value of 0 for $\vartheta$ indicates an exact solution.
@@
    where $d_{1}$ is the projection distance to preference vector, $d_{1}=\frac{{\bm{L}}^{T}{\bm{\lambda}}}{\left\|{\bm{\lambda}}\right\|}$, $d_{2}$ is the distance to ${\bm{L}}$, and $\mu$ is a weight factor, which is set 8.0. The small PBI indicator means the solution is both close to the given preference vector and the true Pareto front.
@@
    The HV indicator measures the quality of a set of objectives $A$.
    $\mathcal{H}_{r}(A)\vcentcolon=\Lambda(\{q\;|\;\exists p\in A:p\preceq q\;\text{and}\;q\preceq r\}),$ (23)
    where $r$ is a reference point and $\Lambda(\cdot)$ serves as the Lebesgue measure.
@@
    The Exact Pareto Optimization (EPO)  [^5] Code: https://github.com/dbmptr/EPOSearch.[14] method. This method to search for exact Pareto solutions before our method. EPO has two steps. The first step is to align the solution with a preference and the second step is to push the solution to the Pareto front. These two steps are repeated until an exact Pareto optimal solution is found. We have fixed numerical bugs in the EPO code. Refer to our code (pgmda.rar) for more details.
@@
    The COSMOS  [^6] Code is from https://github.com/ruchtem/cosmos. [4] method. This method that aims to search for approximate exact Pareto solutions. The COSMOS optimizes the scalarization function,
    $g^{\text{cosmos}}({\bm{\theta}})={\bm{\lambda}}^{T}{\bm{L}}({\bm{\theta}})+\mu S_{c}({\bm{L}}({\bm{\theta}}),{\bm{\lambda}}).$
@@
    Figure 3: Contour of COSMOS.
@@
    $g^{\text{mTche}}({\bm{\theta}})=\max_{i\in[m]}\left\{\frac{{\bm{L}}_{i}(\theta)-\mathbf{z}_{i}}{{\bm{\lambda}}_{i}}\right\},$
    where $z$ is reference point and set as zero vector in this paper. It is known in [15, 37] that under mild conditions, mTche can find the ‘exact’ Pareto solution.
@@ IV-A Synthetic Problems
In this section, we evaluate the performance of PMGDA on six widely-used gradient-based problems: ZDT1[38], MAF1[39], and DTLZ2 [40]. These problems are commonly employed in the gradient-based multi-objective optimization literature [8, 41, 42, 14, 9].
Initial solutions for VLMOP2 are randomly sampled from the range ${[-0.25,0.25]}^{30}$. For the other problems, initial solutions are randomly sampled from ${[0,1]}^{30}$. All methods are run for 500 iterations. For 2-objective problems, 8 preferences are uniformly selected from within the first quadrant, and for 3-objective problems, 15 preferences are uniformly selected from the three-dimensional simplex. The experiments are performed using five different random seeds. $\sigma=0.9$.
@@
TABLE III: Results on all problems
@@
Figure 4: Learning curves on ZDT1.
@@
Figure 5: Learning curves on MAF1.
@@
Figure 6: Learning curves on DTLZ2.
@@
Figure 7: Final exact Pareto solutions on MAF1.
The failure of COSMOS on most problems is primarily attributed to the crucial and inefficient selection of its control parameter setting [4]. As shown in Figure 3,
@@
(a) Circle
(b) Box
Figure 8: Learning curves on MAF1.
@@ IV-B Fairness Classification Task
Figure 9: Result comparison by different methods on the fairness classification problems.
@@
Figure 10: Learning curves on DTLZ2.
@@
Figure 11: ROI constraints for fairness classification problems.
@@ IV-B1 Setting
We test the performance of our method on three well-known fairness classification dataset, namely, Adult [43], Compas [44], and Credit [45]. The first objective ${\bm{L}}_{1}({\bm{\theta}})$ is the cross entropy loss. And the second objective ${\bm{L}}_{2}({\bm{\theta}})$ is the DEO (Difference of Equality of Opportunity) loss [46]. This multiobjective multitask is firstly studied in [4]. During training, batch size is 128. Details of problem are shown in Table IV.
TABLE IV: Fairness classification problem Details.
@@
TABLE V: Results on MTL
@@
 Adult $\vartheta$ 3.54 2.07 2.02 2.19
PBI 0.46 0.41 0.60 0.42
@@ IV-B2 Results
Another advantage of the proposed PMGDA method is its adaptable design of the preference function $h({\bm{\theta}})$. This method is notably the first to successfully identify Pareto solutions that meet specific region of interest (ROI) constraints, as demonstrated in Figure 11. The challenge in applying ROI to multi-task learning problems lies in the stochastic nature of estimated objectives ${\bm{L}}_{i}({\bm{\theta}})$, which vary with batches. Given this estimation noise, we recommend using a smaller, more conservative ROI scale, as applied here (1/1.5).
@@ IV-C Multi-objective Reinforcement Learning (MORL)
Our first finding, illustrated in Figure 12, reveals that our proposed method successfully identifies Pareto optimal solutions within user-defined Regions of Interest (ROIs). Notably, PMGDA is the first gradient-based method capable of meeting such specific demands. In contrast, previous ROI-focused methods relying on evolutionary algorithms, such as those mentioned in [48, 49, 21], have been unable to efficiently optimize policy networks in these tasks.
@@
Figure 13: Exact solutions for the MO-Ant problem, averaged from the final 10 iterations’ evaluations, displayed with error bars.
@@ V Conclusions, Limitations, and Future Works
This paper addresses a fundamental problem in multi-objective optimization: satisfying user-specific preference functions. Unlike previous approaches such as EPO [14], COSMOS [4], and PMTL [8], which rely on specific rules to identify Pareto solutions under preference constraints, our proposed framework is versatile enough to incorporate general preference functions. In the well-known task of finding Pareto solutions that precisely align with a preference vector, our PMGDA method surpasses previous methods like EPO and COSMOS in precision, and outperforms mTche in terms of convergence speed. Additionally, our framework is capable of undertaking new challenges, such as identifying Pareto optimal solutions that meet region of interest constraints—a task rarely addressed in gradient-based Multi-Objective Optimizations (MOOs).

### 16 Feb 2024 v2

@@ IV Experiments
We conducted experiments on three types of problems: synthetic problems, multitask fairness classification problems, and multiobjective reinforcement learning problems. For synthetic problems, the decision vector ${\bm{\theta}}$ is a 30-D vector. For the other two problems, the decision variables are the parameters of neural networks (this number is around 5000-10000).
@@
    The cross angle indicator $\vartheta=\text{arccos}\frac{{\bm{L}}^{\top}{\bm{\lambda}}}{\left\|{\bm{L}}\right\|\left\|{\bm{\lambda}}\right\|}$. A value of 0 for $\vartheta$ indicates an exact solution.
@@
    where $d_{1}$ is the projection distance to preference vector, $d_{1}=\frac{{\bm{L}}^{\top}{\bm{\lambda}}}{\left\|{\bm{\lambda}}\right\|}$, $d_{2}$ is the distance to the objective vector ${\bm{L}}$, and $\mu$ is a weight factor, which is set 8.0. The small PBI indicator means the solution is both close to the given preference vector and the true Pareto front.
@@
    The HV indicator measures the quality of a set of objectives ${\mathbb{A}}$.
    $\mathcal{H}_{\bm{r}}({\mathbb{A}})\vcentcolon=\Lambda(\{{\bm{q}}\;|\;\exists{\bm{p}}\in{\mathbb{A}}:{\bm{p}}\preceq{\bm{q}}\;\text{and}\;{\bm{q}}\preceq{\bm{r}}\}),$ (23)
    where ${\bm{r}}$ is a reference point and $\Lambda(\cdot)$ serves as the Lebesgue measure.
@@
    The Exact Pareto Optimization (EPO)  [^5] Code: https://github.com/dbmptr/EPOSearch.[14] method. This method to search for exact Pareto solutions before our method. EPO has two steps. The first step is to align the solution with a preference and the second step is to push the solution to the Pareto front. These two steps are repeated until an exact Pareto solution is found. We have fixed numerical bugs in the EPO code. Refer to our code (pgmda.rar) for more details.
@@
    The COSMOS  [^6] Code: https://github.com/ruchtem/cosmos. [4] method. This method that aims to search for approximate exact Pareto solutions. The COSMOS optimizes the scalarization function,
    $g^{\text{cosmos}}({\bm{\theta}})={\bm{\lambda}}^{\top}{\bm{L}}({\bm{\theta}})+\mu S_{c}({\bm{L}}({\bm{\theta}}),{\bm{\lambda}}).$
@@
    Figure 3: Contour of COSMOS. For a given preference ${\bm{\lambda}}$, the minimizer of the COSMOS function does not necessary corresponding to the exact Pareto solution (i.e., the COSMOS contour curve is not tangent to the PF).
@@
    $g^{\text{mTche}}({\bm{\theta}})=\max_{i\in[m]}\left\{\frac{L_{i}({\bm{\theta}})-z_{i}}{\lambda_{i}}\right\},$
    where ${\bm{z}}$ is a reference point and is typical set as the ideal point of a MOO problem. It is known in [15, 37] that under mild conditions, mTche can find the ‘exact’ Pareto solution.
@@ IV-A Synthetic Problems
In this section, we evaluate the performance of PMGDA on six widely-used gradient-based problems: ZDT1[38], MAF1[39], and DTLZ2 [40]. These problems are commonly employed in the gradient-based MOO literature [8, 41, 42, 14, 9].
Initial solutions for VLMOP2 are randomly sampled from the range ${[-0.25,0.25]}^{30}$. For the other problems, initial solutions are randomly sampled from ${[0,1]}^{30}$. All methods are run for 500 iterations. For 2-objective problems, 8 preferences are uniformly selected from within the first quadrant, and for 3-objective problems, 15 preferences are uniformly selected from the three-dimensional simplex. The experiments are performed using five different random seeds. The control parameter $\sigma$ in PMGDA is set as 0.9.
@@
TABLE III: Results on synthetic problems. The best solution is marked in bold and the second best is marked with an underline.
@@
Figure 4: Learning curves on ZDT1. The proposed method has the best convergence speed and final performance. The convergence curve of EPO fluctuates before convergence. Final solutions of COSMOS are also not exact Pareto solutions.
@@
Figure 5: Learning curves on MAF1. Final solutions by COSMOS and EPO are not exact Pareto solutions. On this three-objective problem, the convergence of mTche is slow, taking 500 iterations to converge, while PMGDA only needs 100 iterations to converge.
@@
Figure 6: Learning curves on DTLZ2. PMGDA and COSMOS successfully finds all exact Pareto solutions (PMGDA is 3x faster than COSMOS). However, EPO and mTche fail to find all exact Pareto solutions in 500 iterations. The learning curve of EPO is unstable.
@@
Figure 7: Illustrative results on the MAF1 problem. Many solutions searched by EPO and COSMOS are not exact Pareto solutions, i.e., solutions are not aligned with preference vectors.
The failure of COSMOS is attributed to its difficulty of controlling parameter setting [4]. As shown in Figure 3, even this parameter is set as a large value, the optimizer of the COSMOS objective function does not correspond to the exact solution.
@@
(a) Circle constraint
(b) Box constraint
Figure 8: The optimization trajectory in the ROI task. PMGDA successfully find all Pareto solutions satisfying ROI constraints.
@@ IV-B Fairness Classification Task
Figure 9: Result comparison by different methods on the fairness classification problems. Generally, PMGDA solutions align better than other methods. PMGDA find a wider PF than other methods on Adult and Credit and converges better than EPO on Adult. On the most simple Compass problem, EPO and PMGDA perform similiarly.
@@
Figure 10: Result comparisons on Adult and Credit. EPO solutions on Credit is only focused on a small region. PMGDA generally find better exact Pareto solutions on most preferences.
@@
Figure 11: ROI constraints for fairness classification problems. PMGDA find Pareto solutions satisfying the ROI constraint on most preferences. PMGDA is the only method working for ROI constraint.
@@ IV-B1 Setting
We test the performance of our method on three well-known fairness classification dataset, namely, Adult [43], Compas [44], and Credit [45]. The first objective $L_{1}({\bm{\theta}})$ is the cross entropy loss. And the second objective $L_{2}({\bm{\theta}})$ is the DEO (Difference of Equality of Opportunity) loss [46]. This multiobjective multitask is firstly studied in [4]. During training, batch size is 128. The classification model is a three-layer fully-connected model activated by ReLU. Problem and model details are shown in Table IV.
TABLE IV: Details of fairness classification problems.
@@
TABLE V: Results on fairness classification problems. The best result is marked in bold, while the second best result is marked with an underline.
@@
 Adult $\vartheta$downarrow 3.54 2.07 2.02 2.19
PBI $\downarrow$ 0.46 0.41 0.60 0.42
@@ IV-B2 Results
Another advantage of the proposed PMGDA method is its adaptable design of the preference function $h({\bm{\theta}})$. This method is notably the first to successfully identify Pareto solutions that meet specific region of interest (ROI) constraints, as demonstrated in Figure 11. The challenge in applying ROI to multi-task learning problems lies in the stochastic nature of estimated objectives $L_{i}({\bm{\theta}})$, which vary with batches. Given this estimation noise, we recommend using a smaller, more conservative ROI scale, as applied here (1/1.5).
@@ IV-C Multi-objective Reinforcement Learning (MORL)
Our first finding, illustrated in Figure 12, reveals that our proposed method successfully identifies Pareto solutions within user-defined Regions of Interest (ROIs). Notably, PMGDA is the first gradient-based method capable of meeting such specific demands. In contrast, previous ROI-focused methods relying on evolutionary algorithms, such as those mentioned in [48, 49, 21], have been unable to efficiently optimize policy networks in these tasks.
@@
Figure 13: Exact solutions for the MO-Ant problem, averaged from the final 10 iterations’ evaluations, displayed with error bars. PMGDA results outperform mTche results on most subproblems with different preferences.
@@ V Conclusions, Limitations, and Future Works
This paper addresses a fundamental problem in MOO: satisfying user-specific preference functions. Unlike previous approaches such as EPO [14], COSMOS [4], and PMTL [8], which rely on specific rules to identify Pareto solutions under preference constraints, our proposed framework is versatile enough to incorporate general preference functions. In the well-known task of finding Pareto solutions that precisely align with a preference vector, our PMGDA method surpasses previous methods like EPO and COSMOS in precision, and outperforms mTche in terms of convergence speed. Additionally, our framework is capable of undertaking new challenges, such as identifying Pareto solutions that meet region of interest constraints—a task rarely addressed in gradient-based MOOs (MOOs).